OpenStudy (anonymous):

Find the measure of the side indicated with X round to the nearest tenth. Triangle with angles ABC has a 90 degrees at point C and is 41 degrees at point A. From point C to B is 4 cm. What is the length of C to A? Hint : Using tan sign. Thankies :)

OpenStudy (anonymous):



OpenStudy (anonymous):

tan(41 degrees) = 4 / x x = 4 / tan(41 degrees)

OpenStudy (anonymous):

thank you! but what is the final answer for x if we calculate?